# Badminton Birdie-Like Aerodynamic Alignment of Drifting Dust Grains by Subsonic Gaseous Flows in Protoplanetary Disks

### v1

@@ 2.3 Parameterization
While the above discussion allows a qualitative description of alignment, we utilize a parameterization of the double-sphere model to facilitate the quantitative description. In particular, we aim to quantify how $t_{o}$ and $t_{d}$ depend on the characteristic properties of the grain, in addition to the level of asymmetry between the spheres. We define the characteristic length $l$ through
$l\equiv r_{1}+r_{2}$ (18)
@@
$\epsilon\equiv\frac{\varsigma_{1}}{\varsigma_{2}}$ (19)
@@
$\varsigma_{1}+\varsigma_{2}\equiv\lambda l$ (20)
@@
$\varsigma_{1}$ $=\frac{\epsilon\lambda}{1+\epsilon}l,$
$\varsigma_{2}$ $=\frac{\lambda}{1+\epsilon}l.$ (21)
@@
$\kappa\equiv\frac{\rho_{1}}{\rho_{2}}.$ (22)
We denote the total mass of the grain through $m\equiv m_{1}+m_{2}$ and along with definitions of $\epsilon$ and $\kappa$, one can find that
$m_{1}$ $=\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}m,$
$m_{2}$ $=\frac{1}{1+\epsilon^{3}\kappa}m.$ (23)
@@
$r_{1}$ $=\frac{1}{1+\epsilon^{3}\kappa}l,$
$r_{2}$ $=\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}l.$ (24)
@@
$\rho_{1}$ $=\frac{1+\epsilon^{3}}{1+\epsilon^{3}\kappa}\kappa\rho_{s}$
$\rho_{2}$ $=\frac{1+\epsilon^{3}}{1+\epsilon^{3}\kappa}\rho_{s}$ (26)
@@
Figure 3: Dimensionless factors for the double-sphere model as a function of $\epsilon$ (ratio of radii between the two spheres) and $\kappa$ (ratio of material density). Panel a: the damping time dimensionless factor, $\hat{t}_{d,d}$. Panel b: the oscillation time dimensionless factor, $\hat{t}_{o,d}$. As $\epsilon\rightarrow 1$ and $\kappa\rightarrow 1$, $\hat{t}_{o,d}\rightarrow\infty$ meaning there is no oscillation when the two spheres are equivalent.
@@
$t_{d,d}$ $=\frac{\rho_{s}l}{\rho_{g}v_{\text{th}}}\hat{t}_{d,d}\text{ ,}$
$\hat{t}_{d,d}$ $\equiv\frac{2\lambda(1+\epsilon^{3})(1+\epsilon^{3}\kappa)}{\epsilon^{2}(1+\epsilon)(1+\epsilon^{4}\kappa^{2})}\bigg{[}\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}+\frac{2}{5}\lambda^{2}(1+\epsilon^{5}\kappa)\bigg{]}$ (30)
where $\hat{t}_{d,d}$ is a dimensionless factor that encapsulates the asymmetry. We will use the “ $\hat{}$ ” symbol to denote the dimensionless factors that only depend on the geometry of the grain. The physical quantity $\rho_{s}l/(\rho_{g}v_{\text{th}})$ is equivalent to the stopping time of a spherical grain with radius $l$ and material density $\rho_{s}$. The stopping time is the characteristic timescale for the aerodynamic drag force to stop the motion of a moving grain (Armitage, 2015). The relation between the damping time and the stopping time of a single sphere is not too surprising. As the double-sphere grain spins, each sphere moves relative to the gas. Requiring each sphere to stop its relative motion with respect to the gas is equivalent to stopping the spin of the system, i.e., rotational damping. Thus, the damping time should be on the order of the stopping time. However, this effect cannot be captured for a single spherical grain as we will see in Section 3. Plugging in some typical numbers for a protoplanetary disk for the stopping time , we get
@@
Fig. 3a shows the asymmetry factor $\hat{t}_{d,d}$ for different values of size ratio $\epsilon$ and density ratio $\kappa$ while adopting $\lambda=1$ (i.e., contacting spheres). We find that the value is of order unity across the parameter space. When $\epsilon=1$ and $\kappa=1$ (the symmetric case), then $\hat{t}_{d,d}=1.4$. Note that the flow speed $A$ does not contribute to the damping time, but only determines the oscillation potential (and hence frequency) around the dust-gas drift direction.
@@
$t_{o,d}$ $=\sqrt{\frac{\rho_{s}l^{2}}{\rho_{g}v_{\text{th}}A}}\hat{t}_{o,d}\text{ ,}$
$\hat{t}_{o,d}$ $\equiv 2\pi\sqrt{\frac{\lambda(1+\epsilon^{3})}{\epsilon^{2}(1+\epsilon)(\epsilon\kappa-1)}\bigg{[}\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}+\frac{2}{5}\lambda^{2}(1+\epsilon^{5}\kappa)\bigg{]}}$ (32)
where $\hat{t}_{o,d}$ is a dimensionless factor for the oscillation time and the physical quantity in front of it is the characteristic oscillation time. Plugging in the same numbers for a protoplanetary disk as Eq. 2.3 and adopting $A=10$ m/s (see Sec. 4 below), the characteristic oscillation time is $\sim 5$ minutes.
Fig. 3b shows $\hat{t}_{o,d}$ as a function of size ratio $\epsilon$ for different values of density ratio $\kappa$. While most of the parameter space shows $\hat{t}_{o,d}\sim 10$, its value increases rapidly as $\epsilon$ and $\kappa$ decreases to 1. Nevertheless, even if $\kappa=1.01$ and $\epsilon=1$, or $\kappa=1$ and $\epsilon=1.01$, $\hat{t}_{o,d}\sim 50$. Thus, $t_{o,d}$ is much shorter than the damping time and the Keplerian orbital periods of the outer disk even for $1\%$ asymmetry.
@@
$\zeta_{d}$ $=\sqrt{\frac{\rho_{g}v_{\text{th}}}{\rho_{s}A}}\hat{\zeta}_{d}\text{ , }$
$\hat{\zeta}_{d}$ $\equiv\frac{\epsilon(1+\epsilon^{4}\kappa^{2})}{2(1+\epsilon^{3}\kappa)}\sqrt{\frac{(1+\epsilon)^{3}}{\lambda(1+\epsilon^{3})(\epsilon\kappa-1)}}\bigg{[}\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}+\frac{2}{5}\lambda^{2}(1+\epsilon^{5}\kappa)\bigg{]}^{-\frac{1}{2}}$ (33)
where $\hat{\zeta}_{d}$ encapsulates the dimensionless, asymmetry factors. Using typical conditions of disks used in Eq. 2.3, we have $\zeta_{d}\sim 4\times 10^{-7}\hat{\zeta}_{d}$ and $\hat{\zeta}_{d}\sim 6$ for $\epsilon=1.01$ and $\kappa=1$. One can see that $\zeta_{d}\ll 1$ because the gas mass density $\rho_{g}$ is drastically smaller than grain material density $\rho_{s}$ (by 15 orders of magnitude with $n_{g}=10^{9}$ cm-3), which is more than enough to compensate for the difference between the dust-gas relative speed $A$ and the thermal speed $v_{\text{th}}$.
@@
Figure 4: The phase portrait (trajectory map) of the double-sphere grain under the presence of gas flow $\bm{A}$. The horizontal axis is $\theta$, while the vertical axis is the angular velocity $\omega$ normalized by the oscillation frequency $\omega_{o}$. The black circles correspond to stable equilibrium points that serve as the attractors for alignment. The orange plus signs are the unstable equilibrium points. The blue curve is the separatrix that separates the bounded and unbounded regions. Regions inside the separatrix are bounded and oscillate around the attractors.
@@ 3.1 Problem Setup
Assuming specular reflection and subsonic relative motion, Appendix B shows that $d\bm{F}$ at each point of the surface follows
@@ 3.2 Strictly Axisymmteric Spheroid
Thus far, we have not utilized any assumption about the geometry of the spheroid. We now evaluate each term in Eq. 37 using a dedicated coordinate system shown in Fig. 5. The setup is equivalent to Sec. 2. Let $\bm{e}_{1}$, $\bm{e}_{2}$, and $\bm{e}_{3}$ represent the Cartesian, unit directions of the lab frame that is centered at the origin $O$. We define $\bm{e}_{3}$ to be along the direction of gas flow $\bm{A}$ and let the grain rotate around $\bm{e}_{2}$.
Figure 5: Schematic of a prolate and an oblate where the center of mass is shifted along the axis of symmetry. We use ellipses to represent the cross-sections of the two spheroids in the $e_{1}e_{3}$-plane. For the prolate, the axis of symmetry is the long axis, while for the oblate, the axis of symmetry is the short axis. $\bm{e}_{1}$, $\bm{e}_{2}$, and $\bm{e}_{3}$ are lab frame unit vectors where $\bm{e}_{2}$ points out of the page. $\bm{b}_{1}$, $\bm{b}_{2}$, and $\bm{b}_{3}$ are body frame unit vectors, where $\bm{b}_{2}$ also points out of the page and $\bm{b}_{3}$ is along the axis of symmetry. The offset of the center of the spheroid from the center of mass $O$ is $\bm{g}$. Since the center of mass is shifted along the axis of symmetry, $\bm{g}\parallel\bm{b}_{3}$. The green arrow is the direction of the gas flow $\bm{A}$.
Let the body frame of the spheroid also be a Cartesian coordinate system with its center denoted by $G$. We use $\bm{b}_{1}$, $\bm{b}_{2}$, and $\bm{b}_{3}$ to denote the unit vectors. $\bm{b}_{3}$ is along the axis of asymmetry and $\bm{b}_{2}$ is parallel to $\bm{e}_{2}$. The angle between $\bm{b}_{3}$ and $\bm{e}_{3}$ is $\theta$. The relation between these two sets of unit vectors is simply
$\bm{b}_{i}=W_{ij}\bm{e}_{j}$ (38)
@@
$\frac{x^{2}+y^{2}}{a^{2}}+\frac{z^{2}}{c^{2}}=1$ (40)
where $x$, $y$, and $z$ are coordinates in the body frame along $\bm{b}_{1}$, $\bm{b}_{2}$, and $\bm{b}_{3}$, respectively. In addition, $c$ is the length of the spheroid along the axis of symmetry (not to be confused with the isothermal sound speed $c_{s}$ or the speed of light) and $a$ is the length along the axis perpendicular to that. A prolate has $c>a$, while an oblate has $c<a$.
Considering axisymmetric particles makes the surface integral easier to compute via a surface of revolution. We parameterize the surface by introducing the azimuthal angle $\phi$ around the $\bm{b}_{3}$ axis
$\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}w(z)\cos\phi\\ w(z)\sin\phi\end{pmatrix}$ (41)
where $\phi$ starts from $\bm{b}_{1}$ and increases following the right-hand rule. Furthermore, $w(z)$ is the perpendicular distance from the surface to the axis of symmetry and serves as the generatrix for the surface integration:
$w(z)\equiv a\sqrt{1-\bar{z}^{2}}$ (42)
where $\bar{z}\equiv z/c$ is $z$ normalized by $c$. We will use the bar symbol to denote length quantities normalized by $c$. With the generatrix, we can derive the surface element to be
$d\sigma=ac\sqrt{1+(\bar{a}^{2}-1)\bar{z}^{2}}d\bar{z}d\phi$ (43)
@@
The normal vector of the spheroid is related to the gradient of Eq. 40. We express the unit normal vector by $\bm{n}=\tilde{n}_{i}\bm{b}_{i}$:
$\begin{pmatrix}\tilde{n}_{1}\\ \tilde{n}_{2}\\ \tilde{n}_{3}\end{pmatrix}=\frac{1}{\sqrt{1+(\bar{a}^{2}-1)\bar{z}^{2}}}\begin{pmatrix}\sqrt{1-\bar{z}^{2}}\cos\phi\\ \sqrt{1-\bar{z}^{2}}\sin\phi\\ \bar{a}\bar{z}\end{pmatrix}\text{ .}$ (44)
We use a tilde to denote components using $\bm{b}_{i}$ as basis vectors.
@@
where $\bm{g}$ points from the center of mass to the geometric center $G$ and $\bm{s}$ points from $G$ to the surface. The components of $\bm{s}$ using $\bm{b}_{i}$ basis vectors are
$\begin{pmatrix}\tilde{s}_{1}\\ \tilde{s}_{2}\\ \tilde{s}_{3}\end{pmatrix}=\begin{pmatrix}a\sqrt{1-\bar{z}^{2}}\cos\phi\\ a\sqrt{1-\bar{z}^{2}}\sin\phi\\ c\bar{z}\end{pmatrix}$ (46)
For the offset vector $\bm{g}$, we first consider the offset along the axis of symmetry to follow the strict assumption of axisymmetry. If the length between $O$ and $G$ is $g$, we have $\bm{g}=g\bm{b}_{3}$. We later relax this assumption and explore the effects of $\bm{g}$ when it is not along the axis of symmetry.
@@
where $f(\bar{z})$ is some function that depends on $\bar{z}$. Although $\bar{z}$ serves as a dummy variable for integration under this definition, it originates from the coordinate along the symmetry axis of the spheroid as we integrate over the surface. The definition of $E[f(\bar{z})]$ implicitly depends on $\bar{a}$ for easier notation. Appendix C shows the analytical solutions to the $E$-integrals used in this paper.
@@
It is convenient to know that the integral of N expressed in $\bm{b}_{i}$ basis vectors is
@@
For the first term of Eq. 3.2, using the $\bm{b}_{i}$ basis vectors, we find that
$\bm{g}\times\oint_{S}\textbf{{N}}d\sigma(\bm{g}\times)=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-\bar{z}^{2}]\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}$ (51)
@@
$\oint_{S}\bm{s}\times\textbf{{N}}(\bm{s}\times)d\sigma=-\pi c^{4}\bar{a}(1-\bar{a}^{2})^{2}E[\bar{z}^{2}-\bar{z}^{4}]\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}$ (52)
@@
$\tilde{L}_{11}=\tilde{L}_{22}=-\pi c^{4}\bar{a}\bigg{[}\bar{g}^{2}E[1-\bar{z}^{2}]+(1-\bar{a}^{2})^{2}E[\bar{z}^{2}-\bar{z}^{4}]\bigg{]}$ (53)
@@
Lastly, for M, using the $\bm{b}_{i}$ basis vectors, we have
@@
$=\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$ (54)
@@
We can now derive the torque from Eq. 37. Given the assumed $\bm{A}=A\bm{e}_{3}$, we can express $\bm{A}=\tilde{A}_{i}\bm{b}_{i}$ and find that
$\bm{A}=-A\sin\theta\bm{b}_{1}+A\cos\theta\bm{b}_{3}$ (55)
using Eq. 39. Since we only consider rotation around $\bm{e}_{2}$ and $\bm{e}_{2}=\bm{b}_{2}$, the angular velocity is simply $\bm{\omega}=\omega\bm{e}_{3}=\omega\bm{b}_{2}$. Applying to Eq. 37, the torque is thus
$\tilde{\Gamma}_{2}=$ $-\rho_{g}v_{\text{th}}\pi c^{4}\omega\bar{a}\bigg{[}\bar{g}^{2}E[1-\bar{z}^{2}]+(1-\bar{a}^{2})^{2}E[\bar{z}^{2}-\bar{z}^{4}]\bigg{]}$
$-\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]\sin\theta$ (56)
and we find that $\tilde{\Gamma}_{1}=\tilde{\Gamma}_{3}=0$. Note that $\Gamma_{2}=\tilde{\Gamma}_{2}$ also because $\bm{e}_{2}=\bm{b}_{2}$. Thus, we see that the torque is only limited along $\bm{e}_{2}$ which is what we expect when we only consider $\bm{\omega}$ along $\bm{e}_{2}$.

### v2

@@ 2.3 Parameterization
While the above discussion allows a qualitative description of alignment, we utilize a parameterization of the double-sphere model to facilitate the quantitative description. In particular, we aim to quantify how $t_{o}$ and $t_{d}$ depend on the characteristic properties of the grain, in addition to the level of asymmetry between the spheres. We define the characteristic length $l$ of the entire grain through
$l\equiv r_{\bullet}+r_{\circ}$ (18)
@@
$\epsilon\equiv\frac{\varsigma_{\bullet}}{\varsigma_{\circ}}$ (19)
@@
$\varsigma_{\bullet}+\varsigma_{\circ}\equiv\lambda l$ (20)
@@
$\varsigma_{\bullet}$ $=\frac{\epsilon\lambda}{1+\epsilon}l,$
$\varsigma_{\circ}$ $=\frac{\lambda}{1+\epsilon}l.$ (21)
@@
$\kappa\equiv\frac{\rho_{\bullet}}{\rho_{\circ}}.$ (22)
We denote the total mass of the grain through $m\equiv m_{\bullet}+m_{\circ}$ and along with definitions of $\epsilon$ and $\kappa$, one can find that
$m_{\bullet}$ $=\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}m,$
$m_{\circ}$ $=\frac{1}{1+\epsilon^{3}\kappa}m.$ (23)
@@
$r_{\bullet}$ $=\frac{1}{1+\epsilon^{3}\kappa}l,$
$r_{\circ}$ $=\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}l.$ (24)
@@
$\rho_{\bullet}$ $=\frac{1+\epsilon^{3}}{1+\epsilon^{3}\kappa}\kappa\rho_{s}$
$\rho_{\circ}$ $=\frac{1+\epsilon^{3}}{1+\epsilon^{3}\kappa}\rho_{s}$ (26)
@@
Figure 3: Dimensionless factors for the double-sphere model as a function of $\epsilon$ (ratio of radii between the two spheres) and $\kappa$ (ratio of material density). Panel a: the damping time dimensionless factor, $\breve{t}_{d,d}$. Panel b: the oscillation time dimensionless factor, $\breve{t}_{o,d}$. As $\epsilon\rightarrow 1$ and $\kappa\rightarrow 1$, $\breve{t}_{o,d}\rightarrow\infty$ meaning there is no oscillation when the two spheres are equivalent.
@@
$t_{d,d}$ $=\frac{\rho_{s}l}{\rho_{g}v_{\text{th}}}\breve{t}_{d,d}\text{ ,}$
$\breve{t}_{d,d}$ $\equiv\frac{2\lambda(1+\epsilon^{3})(1+\epsilon^{3}\kappa)}{\epsilon^{2}(1+\epsilon)(1+\epsilon^{4}\kappa^{2})}\bigg{[}\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}+\frac{2}{5}\lambda^{2}(1+\epsilon^{5}\kappa)\bigg{]}$ (30)
where $\breve{t}_{d,d}$ is a dimensionless factor that encapsulates the asymmetry. We will use a breve ( $\breve{}$ ) to denote the dimensionless factors that only depend on the geometry of the grain. The physical quantity $\rho_{s}l/(\rho_{g}v_{\text{th}})$ is equivalent to the stopping time of a spherical grain with radius $l$ and material density $\rho_{s}$. The stopping time is the characteristic timescale for the aerodynamic drag force to stop the motion of a moving grain (Armitage, 2015). The relation between the damping time and the stopping time of a single sphere is not too surprising. As the double-sphere grain spins, each sphere moves relative to the gas. Requiring each sphere to stop its relative motion with respect to the gas is equivalent to stopping the spin of the system, i.e., rotational damping. Thus, the damping time should be on the order of the stopping time. However, this effect cannot be captured for a single spherical grain as we will see in Section 3. Plugging in some typical numbers for a protoplanetary disk for the stopping time, we get
@@
Fig. 3a shows the asymmetry factor $\breve{t}_{d,d}$ for different values of size ratio $\epsilon$ and density ratio $\kappa$ while adopting $\lambda=1$ (i.e., contacting spheres). We find that the value is of order unity across the parameter space. When $\epsilon=1$ and $\kappa=1$ (the symmetric case), then $\breve{t}_{d,d}=1.4$. Note that the flow speed $A$ does not contribute to the damping time, but only determines the oscillation potential (and hence frequency) around the dust-gas drift direction.
@@
$t_{o,d}$ $=\sqrt{\frac{\rho_{s}l^{2}}{\rho_{g}v_{\text{th}}A}}\breve{t}_{o,d}\text{ ,}$
$\breve{t}_{o,d}$ $\equiv 2\pi\sqrt{\frac{\lambda(1+\epsilon^{3})}{\epsilon^{2}(1+\epsilon)(\epsilon\kappa-1)}\bigg{[}\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}+\frac{2}{5}\lambda^{2}(1+\epsilon^{5}\kappa)\bigg{]}}$ (32)
where $\breve{t}_{o,d}$ is a dimensionless factor for the oscillation time and the physical quantity in front of it is the characteristic oscillation time. Plugging in the same numbers for a protoplanetary disk as Eq. 2.3 and adopting $A=10$ m/s (see Sec. 4 below), the characteristic oscillation time is $\sim 5$ minutes.
Fig. 3b shows $\breve{t}_{o,d}$ as a function of size ratio $\epsilon$ for different values of density ratio $\kappa$. While most of the parameter space shows $\breve{t}_{o,d}\sim 10$, its value increases rapidly as $\epsilon$ and $\kappa$ decreases to 1. Nevertheless, even if $\kappa=1.01$ and $\epsilon=1$, or $\kappa=1$ and $\epsilon=1.01$, $\breve{t}_{o,d}\sim 50$. Thus, $t_{o,d}$ is much shorter than the damping time and the Keplerian orbital periods of the outer disk even for $1\%$ asymmetry.
@@
$\zeta_{d}$ $=\sqrt{\frac{\rho_{g}v_{\text{th}}}{\rho_{s}A}}\breve{\zeta}_{d}\text{ , }$
$\breve{\zeta}_{d}$ $\equiv\frac{\epsilon(1+\epsilon^{4}\kappa^{2})}{2(1+\epsilon^{3}\kappa)}\sqrt{\frac{(1+\epsilon)^{3}}{\lambda(1+\epsilon^{3})(\epsilon\kappa-1)}}\bigg{[}\frac{\epsilon^{3}\kappa}{1+\epsilon^{3}\kappa}+\frac{2}{5}\lambda^{2}(1+\epsilon^{5}\kappa)\bigg{]}^{-\frac{1}{2}}$ (33)
where $\breve{\zeta}_{d}$ encapsulates the dimensionless, asymmetry factors. Using typical conditions of disks used in Eq. 2.3, we have $\zeta_{d}\sim 4\times 10^{-7}\breve{\zeta}_{d}$ and $\breve{\zeta}_{d}\sim 6$ for $\epsilon=1.01$ and $\kappa=1$. One can see that $\zeta_{d}\ll 1$ because the gas mass density $\rho_{g}$ is drastically smaller than grain material density $\rho_{s}$ (by 15 orders of magnitude with $n_{g}=10^{9}$ cm-3), which is more than enough to compensate for the difference between the dust-gas relative speed $A$ and the thermal speed $v_{\text{th}}$.
@@
Figure 4: The phase portrait (trajectory map) of the double-sphere grain under the presence of gas flow $\bm{A}$. The horizontal axis is $\theta$, while the vertical axis is the angular velocity $\omega$ normalized by the oscillation frequency $\omega_{o}$. The black circles correspond to stable equilibrium points that serve as the attractors for alignment. The plus signs are the unstable equilibrium points. The blue curve is the separatrix that separates the bounded and unbounded regions. Regions inside the separatrix are bounded and oscillate around the attractors.
@@ 3.1 Problem Setup
Assuming specular reflection, gas drag in the Epstein regime, and subsonic relative motion, Appendix B shows that $d\bm{F}$ at each point of the surface follows
@@ 3.2 Strictly Axisymmteric Spheroid
Thus far, we have not utilized any assumption about the geometry of the spheroid. We now evaluate each term in Eq. 37 using a dedicated coordinate system shown in Fig. 5. The setup is equivalent to Sec. 2. Let $\bm{\hat{e}}_{1}$, $\bm{\hat{e}}_{2}$, and $\bm{\hat{e}}_{3}$ represent the Cartesian, unit directions of the lab frame that is centered at the origin $O$. We define $\bm{\hat{e}}_{3}$ to be along the direction of gas flow $\bm{A}$ and let the grain rotate around $\bm{\hat{e}}_{2}$.
Figure 5: Schematic of a prolate and an oblate where the center of mass is shifted along the axis of symmetry. We use ellipses to represent the cross-sections of the two spheroids in the $e_{1}e_{3}$-plane. The quantity $a$ is the length from the spheroid center along the axis perpendicular to the axis of symmetry, while $c$ is the length from the center along the axis of symmetry. For the prolate, the axis of symmetry is the long axis ($c>a$), while for the oblate, the axis of symmetry is the short axis ($c<a$). $\bm{\hat{e}}_{1}$, $\bm{\hat{e}}_{2}$, and $\bm{\hat{e}}_{3}$ are lab frame unit vectors where $\bm{\hat{e}}_{2}$ points out of the page. $\bm{\hat{b}}_{1}$, $\bm{\hat{b}}_{2}$, and $\bm{\hat{b}}_{3}$ are body frame unit vectors, where $\bm{\hat{b}}_{2}$ also points out of the page and $\bm{\hat{b}}_{3}$ is along the axis of symmetry. The offset of the center of the spheroid from the center of mass $O$ is $\bm{g}$. Since the center of mass is shifted along the axis of symmetry, $\bm{g}\parallel\bm{\hat{b}}_{3}$. The green arrow is the direction of the gas flow $\bm{A}$.
Let the body frame of the spheroid also be a Cartesian coordinate system with its center denoted by $G$. We use $\bm{\hat{b}}_{1}$, $\bm{\hat{b}}_{2}$, and $\bm{\hat{b}}_{3}$ to denote the unit vectors. $\bm{\hat{b}}_{3}$ is along the axis of asymmetry and $\bm{\hat{b}}_{2}$ is parallel to $\bm{\hat{e}}_{2}$. The angle between $\bm{\hat{b}}_{3}$ and $\bm{\hat{e}}_{3}$ is $\theta$. The relation between these two sets of unit vectors is simply
$\bm{\hat{b}}_{i}=W_{ij}\bm{\hat{e}}_{j}$ (38)
@@
$\frac{b_{1}^{2}+b_{2}^{2}}{a^{2}}+\frac{b_{3}^{2}}{c^{2}}=1$ (40)
where $b_{1}$, $b_{2}$, and $b_{3}$ are coordinates in the body frame along $\bm{\hat{b}}_{1}$, $\bm{\hat{b}}_{2}$, and $\bm{\hat{b}}_{3}$, respectively. In addition, $c$ is the length of the spheroid along the axis of symmetry (not to be confused with the isothermal sound speed $c_{s}$ or the speed of light) and $a$ is the length along the axis perpendicular to that. A prolate has $c>a$, while an oblate has $c<a$. Note that to fulfill the assumption of drag in the Epstein regime, the longest axis of the spheroid should be smaller than the mean free path of the gas (which is estimated in Sec. 4.2 below for an illustrative disk).
Considering axisymmetric particles makes the surface integral easier to compute via a surface of revolution. We parameterize the surface by introducing the azimuthal angle $\phi$ around the $\bm{\hat{b}}_{3}$ axis
$\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}=\begin{pmatrix}w\cos\phi\\ w\sin\phi\end{pmatrix}$ (41)
where $\phi$ starts from $\bm{\hat{b}}_{1}$ and increases following the right-hand rule. Furthermore, $w$ is the perpendicular distance from the surface to the axis of symmetry, which depends on $b_{3}$, and serves as the generatrix for the surface integration:
$w(b_{3})\equiv a\sqrt{1-\bar{b}_{3}^{2}}$ (42)
where $\bar{b}_{3}\equiv b_{3}/c$ is $b_{3}$ normalized by $c$. We will use the bar symbol to denote length quantities normalized by $c$. With the generatrix, we can derive the surface element to be
$d\sigma=ac\sqrt{1+(\bar{a}^{2}-1)\bar{b}_{3}^{2}}d\bar{b}_{3}d\phi$ (43)
@@
The normal vector of the spheroid is related to the gradient of Eq. 40. We express the unit normal vector by $\bm{n}=\tilde{n}_{i}\bm{\hat{b}}_{i}$:
$\begin{pmatrix}\tilde{n}_{1}\\ \tilde{n}_{2}\\ \tilde{n}_{3}\end{pmatrix}=\frac{1}{\sqrt{1+(\bar{a}^{2}-1)\bar{b}_{3}^{2}}}\begin{pmatrix}\sqrt{1-\bar{b}_{3}^{2}}\cos\phi\\ \sqrt{1-\bar{b}_{3}^{2}}\sin\phi\\ \bar{a}\bar{b}_{3}\end{pmatrix}\text{ .}$ (44)
We use a tilde to denote components using $\bm{\hat{b}}_{i}$ as basis vectors.
@@
where $\bm{g}$ points from the center of mass to the geometric center $G$ and $\bm{s}$ points from $G$ to the surface. The components of $\bm{s}$ using $\bm{\hat{b}}_{i}$ basis vectors are
$\begin{pmatrix}\tilde{s}_{1}\\ \tilde{s}_{2}\\ \tilde{s}_{3}\end{pmatrix}=\begin{pmatrix}a\sqrt{1-\bar{b}_{3}^{2}}\cos\phi\\ a\sqrt{1-\bar{b}_{3}^{2}}\sin\phi\\ c\bar{b}_{3}\end{pmatrix}$ (46)
For the offset vector $\bm{g}$, we first consider the offset along the axis of symmetry to follow the strict assumption of axisymmetry. If the length between $O$ and $G$ is $g$, we have $\bm{g}=g\bm{\hat{b}}_{3}$. We later relax this assumption and explore the effects of $\bm{g}$ when it is not along the axis of symmetry.
@@
where $x$ is a dummy variable for integration and $f(x)$ is some function that depends on $x$. The dummy variable $x$ originates from $\bar{b}_{3}$ as we integrate over the surface. The definition of $E[f(x)]$ implicitly depends on $\bar{a}$ for easier notation. Appendix C shows the analytical solutions to the $E$-integrals used in this paper.
@@
It is convenient to know that the integral of N expressed in $\bm{\hat{b}}_{i}$ basis vectors is
@@
For the first term of Eq. 3.2, using the $\bm{\hat{b}}_{i}$ basis vectors, we find that
$\bm{g}\times\oint_{S}\textbf{{N}}d\sigma(\bm{g}\times)=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-x^{2}]\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}$ (51)
@@
$\oint_{S}\bm{s}\times\textbf{{N}}(\bm{s}\times)d\sigma=-\pi c^{4}\bar{a}(1-\bar{a}^{2})^{2}E[x^{2}-x^{4}]\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{pmatrix}$ (52)
@@
$\tilde{L}_{11}=\tilde{L}_{22}=-\pi c^{4}\bar{a}\bigg{[}\bar{g}^{2}E[1-x^{2}]+(1-\bar{a}^{2})^{2}E[x^{2}-x^{4}]\bigg{]}$ (53)
@@
Lastly, for M, using the $\bm{\hat{b}}_{i}$ basis vectors, we have
@@
$=\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$ (54)
@@
We can now derive the torque from Eq. 37. Given the assumed $\bm{A}=A\bm{\hat{e}}_{3}$, we can express $\bm{A}=\tilde{A}_{i}\bm{\hat{b}}_{i}$ and find that
$\bm{A}=-A\sin\theta\bm{\hat{b}}_{1}+A\cos\theta\bm{\hat{b}}_{3}$ (55)
using Eq. 39. Since we only consider rotation around $\bm{\hat{e}}_{2}$ and $\bm{\hat{e}}_{2}=\bm{\hat{b}}_{2}$, the angular velocity is simply $\bm{\omega}=\omega\bm{\hat{e}}_{3}=\omega\bm{\hat{b}}_{2}$. Applying to Eq. 37, the torque is thus
$\tilde{\Gamma}_{2}=$ $-\rho_{g}v_{\text{th}}\pi c^{4}\omega\bar{a}\bigg{[}\bar{g}^{2}E[1-x^{2}]+(1-\bar{a}^{2})^{2}E[x^{2}-x^{4}]\bigg{]}$
$-\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]\sin\theta$ (56)
and we find that $\tilde{\Gamma}_{1}=\tilde{\Gamma}_{3}=0$. Note that $\Gamma_{2}=\tilde{\Gamma}_{2}$ also because $\bm{\hat{e}}_{2}=\bm{\hat{b}}_{2}$. Thus, we see that the torque is only limited along $\bm{\hat{e}}_{2}$ which is what we expect when we only consider $\bm{\omega}$ along $\bm{\hat{e}}_{2}$.